A Complete Bibliography of Publications in Applied Mathematics and Computation: 1985–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 December 2023 Version 1.05

Title word cross-reference

2 [Zak89]. 3 [TT86, TT88]. b [KKS87]. C^{1} [Sai88]. h [KKS87]. k [BR88]. m [MS87b]. M(t)/G/∞ [Pat87, PM88]. Q [HDP85]. RL [LCE85]. u_{xyz} = f [DR88]. V [Par85]. W [WH85]. x′ = y, y′ = 0 [RK87]. Z [VA88].

-analysis [HDP85]. -cycle [Par85]. -dimensional [BR88]. -homogeneous [MS87b]. -transform [VA88].

20 [LMM86]. 200 [LMM86]. 24 [LMM86].

34 [Ano89a].

abelian [Pig87]. Absorbing [Soc88]. accelerating [GW89]. action [IA89]. actions [Mat86]. adaptive [And87, HMOT86, NN87, Shy87]. advanced [HMOT86]. aesthetic [Mur89a, Mur89b]. afterbody [Hof88]. AI [Low86]. Airflow [MKJ87]. airfoil [Gol88a]. algebra [DS86]. Algebraic [Bra86, Hof88, Man88, Gre87b, HT88, KM87]. algorithm [CK86, Ge89c, LE85, LCE85, Rav87, WC88]. algorithms [CS85, HMOT86, HT88, KM86, MHL86, Mel86, Par89, TT88]. allocation [AB87, Les89]. allowable [Jam85]. alternating [CG88]. AMG [Rug86]. analyses [PM88]. Analysis [Axe86, Sas85, EBR85, HDP85, HC87a, Jam85, KR89, KL89, LCE85, pLR87a, pLR87b, Mat86, Pal86, SR88, VJ88, Zel87, ZA88]. Analytic [Ado88, Har87, IA85, Mil88a, SV85]. Analytical [AP87, IA85]. anomalies [ZA88]. aperiodic [Gae87]. Application [AB87, Ioa88, Mer86, RS86, WC88, AP87, Har87, KKS87, RMR86, Sos88]. Applications [GS86]. applied [Ger85, Kam88, Ano89a]. Applying [MCD89]. approach [FMR86, GH89, Gre87a, IA85, JB89a, Uen89]. Approximate [KR87a, Vál88]. approximation [CK86]. approximations [MKJ87]. architectures [Pop86]. areas [DG87a]. argument [AW87]. arguments [KR87b]. arid [Fis85]. arising [SW89]. ARMA [NH86]. Aspects [KW86, NS89]. asset [Ree87]. assignable [JB89b]. associated [Ber89]. astrophysical [BCZ89]. Asymptotic [KR89, LP89, RSS89, Ado87, Boy89]. attitude [HH85]. audit [Ais88]. Author [Ano86a]. autocorrelated [HC87b]. automata [Hog88]. Automatic [Wex87, Ott86]. Automation [KPT87]. averaging [BM89]. avoid [Mor86b].

balance [Pat89]. banded [Sca89]. based [IA89, Sug87]. bases [Yse86]. Basic [TT86]. basis [Boy89]. Bayesian [FMR86]. behavior [Gae87, RSS89, SV85]. Bergman [RMR86]. Bergman-type [RMR86]. best [HQU88]. between [Fra89, Jam85, Ott86]. Bifurcations [CS88, MR89]. Bioeconomic [Mos88b]. Biology [Con89, CS89, Les89]. Black [Den86, Den88, KL89]. blades [WC88]. board [Ano85, Ano86b, Ano88b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88c, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f]. Boolean [TTB88]. borehole [ZA88]. Boundary [Zam88, AG85, Aga86, HL88b, HL88a, Ier86, KR87c, Mit86, Oma86, Pan88, SW89, Soc88, UA88, WKTM88]. boundary-value [WKTM88]. bounded [Sha87, Sos88]. bounding [Wol89]. bounds [AG89]. box [Den86, Den88]. bundle [Hof89]. calculation [GL86, LCE85]. calculations [Mel86]. calculus [TTB88]. canonical [IA89]. case [Bra86, Got85, Ier86]. catastrophes [MR89]. cathodic [Zam88]. Cauchy [Gol88b, Gol89, Uen89]. causality [NH86]. causes [JB89b]. Cellular [Hog88]. cement [Ger85]. cement-plant [Ger85]. chain [Ben89]. changes [Mos88a, Tur88]. changing [VB89]. charge [VJ88]. charts [JB89a, JB89b]. chasing [AG85]. Chebyshev [Boy89, Lon87c]. chemical [Shi89]. chemostat [SW89]. choice [Ge89a, GFM88, Gil89]. Circular [Car88a]. class [AG89, Ge89b, GW89, IP89]. closed [IP89]. closed-form [IP89]. Coefficient [MS89, RS86]. Coefficient-parameter [MS89]. coefficients [Boy89, Gol88b, Gol89]. Collocation [Hei88]. color [KL89]. commutation [Mur89b]. Comparison [Sca89, CS85]. comparisons [HQU88, Mat86]. compartmental [Pat89]. compartmentalized [PM89]. competition [Les89, SV85, SW89]. compiler [KN86]. compiler-driven [KN86]. Complementary [Lon87a]. complex [DLR85, Mei89, MW88]. complexity [BR88]. composite [HMOT86]. composite-grid [HMOT86]. Computation [Ano89a, DLR85, LE85, Shy87, SJK86]. Computational [Aga86, Kun86, CAP88, Hic86, KKS87, MHL86]. Computations [Lar86]. Computer [HC87a, TW89, Gre87a, Pop86]. computer-oriented [Gre87a]. computers [Axe86, DS86, HMOT86]. Computing [MS87a, Ge89b]. concepts [HDP85, Mer86, WHF89]. concordance [HDP85]. condition [MM85, Shi89]. conditions [Beh89, Soc88]. conflicts [Szi89]. conjugate [Kap88, Yse86]. conjugate-gradient [Kap88]. conservative [VJ88]. considerations [Kun86, pLR87a]. constant [AW87, Gol88b, Gol89]. constants [DT87]. constrained [KM86]. constructing [Oma86]. construction [Eis87, WC88]. consumer [Gae87]. contact [Hof89]. continuation [MMP89b, Mit86, MS87a, MS89]. continuous [GH89, PM89]. continuum [Hic86, MHL86]. contour [WC88]. contractions [ZA88]. control [BM89, BdA89, CK86, Fei85, HH85, JB89a, JB89b]. controls [HH85]. convergence [GW89, HK89, KM87, KH87, Lon87b, pLR87a, MM87, Man86, Par85, Par87, Yse86]. convex [Wol89]. coordinates [Bar88]. cortex [Hof89]. cosine [EE89]. countries [Zem86]. county [Tur88]. coupling [Mei89]. Cray [LMM86]. criterion [Mat86]. cubic [Beh89]. cumulative [Pat89]. curves [Eis87, Ioa88]. cyber [Sca89]. cycle [Par85, TT89]. Cyclic [Ais88].

D [TT86, TT88, Zak89]. damping [Whi89]. data [Ben89, EBR85, Mer86, MKJ87, Pat89]. Decision [DDS85, BDT86, MCD89, Rav87]. decomposition [Ado87]. decompositions [CS88]. defined [Ioa88, Kee89]. definite [EE89, Man88]. degenerate [RK87]. delay [AW87]. dependence [Mit86]. dependent [ZH86]. Derivation [Fis85, IA85]. derivative [KPT87]. Derivatives [Har87, Wex87]. derived [Beh89]. design [CS85, JB89b, Sug87]. detection [Ott86]. Determinacy [RK87]. determination [BJ85, Ioa88]. determining [SR88]. deterministic [GD87]. developed [Zem86]. development [Mei89, Nic86]. deviating [KR87b]. Diagonalization [LR85]. diced [Lan89]. difference [GP88, Pan88, Sas88, ZA88]. different [Mur89a]. Differential [AW87, Ada89, AR87, DT87, LS87, LL89, LP89, Mik87, MB88, Nie85, SR88]. differentiation [MM85]. diffusion [AG89, Mei89]. dilemma [NS89]. dimensional [BR88, Hol86, KM87, ZA88, Zia89]. dimensions [Arn86, CG87, Kap88, KW86]. dipole [VJ88]. direct [KB85, bS89]. directly [bS89]. discordance [HDP85]. Discrete [Gol89, IA89, Aga86, DS87, GS86, Got85, PM89, Sha87, UA88, VJ86, WH85]. disease [Wat89]. disequilibrium [Zha89]. disk [Bar88]. displacement [KKS87]. distance [Fra89]. distributed [BW88]. distributions [Eis87, PM89]. domain [ZA88]. dominance [Kee89]. doubling [AB87]. Drilling [DG87a]. driven [KN86]. duality [Vál88]. Due [DDS85]. during [Mei89]. dynamic [Hen86, Udw87a, Udw87b]. dynamical [BL87, MR89, NS89]. dynamics [CJ88, DS87, Hic86, Jam88, MHL86].

earning [Ree87]. Eckalbar [Zha89]. ecological [Hog88]. econometric [MP86, Pal86]. econometrics [Hen86]. economy [Sos88]. Editorial [Ano85, Ano86b, Ano88b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88c, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f]. effect [TW89]. effective [JB89a]. Efficient [KM86, Ada89]. efficiently [NH86]. elastic [Soc88]. elasticity [KM87, Rug86]. electrochemistry [LE85]. electron [Kun86]. element [CG87, Cha88, Zam88]. elementary [IP89]. elements [Pig87]. eligible [Pig87]. elliptic [Ada89, Bar89, CG87, GP88, KB85, RS86, TT86]. ELXSI [SM88]. Empirical [Hen86]. End [Beh89]. endpoint [Boy89]. energy [Got85]. Enumerating [Pig87]. environment [Nic86, OM86, VB89]. environmental [Jam86]. epidemics [Gle88]. epidemiological [TW89]. equation [AP87, BdA89, Bar88, BYPRV86, DR88, Gol88a, HL86, Hol86, HL88a, JPS86, RS86, Sug87]. equations [Ada89, AR87, Ado87, Ado88, AW87, AP87, Bee87, Cha88, DT87, Gol88b, Gol89, Got85, HL88b, IP89, IA89, KB85, Kad86, KKS87, LV87, LS87, Lam86, MMP89a, Mer86, Mik87, Mil87, MB88, NN87, Nie85, RS86, Sas85, Sas88, Sca89, Sha89, SR88, Soc88, ZH86, Zia89]. Equidistribution [And87]. equilibria [RK87]. equipotential [Ioa88]. equivalence [GP88]. equivalents [GD87]. Error [CG88, Ber89, LCE85]. estimate [Pet88]. estimates [CG88, KM87, MM87, Pat87, Pat89]. estimating [GFM88, Gil89]. Estimation [Whi89, Zia89, CSR88, MW88]. Euler [Mer86, SJK86]. eutrophication [Mos88b]. Evaluation [EE89, LMM86, Wex87]. evolutionary [VB89, CV89]. exact [Fei85, IA85]. example [Ber89]. excited [Gre87b]. Existence [Sas88]. expansion [RMR86]. Expected [pLR87a, pLR87b]. experimental [Ben89, EBR85, pLR87a]. experiments [HC87a, SM88]. Explicit [AR87, HH85]. exponential [MMP89a]. exponentially [HC87b]. expression [Ott86]. extension [Ge89a, GW89, Smi86]. extremal [Oma86]. Extreme [HC87b, HC87a]. extreme-value [HC87a].

FAC [HMOT86]. factorable [Ge89c]. factorizations [Axe86, Kap88]. FAS [RS86]. fast [Bar89, HMOT86, KB85]. fatal [Wat89]. feedback [BL87, HH85]. field [Mur89a, Mur89b, VJ88]. filariasis [TT89]. filtering [Cha89]. Finding [Mil87, Mil88a]. Finite [Car88b, Cha88, ZA88, CG87, Hof88]. Finite-difference [ZA88]. finite-span [Hof88]. fins [Hof88]. first [CG87, LS87, Nie85]. flow [Mel86, Mer86, Pat89, Shy87]. flows [SJK86, ZA88]. fluids [BS86]. forecasting [EB86]. form [IP89]. formation [Mei89]. forms [BR88]. formula [DT87, Gre87b, Lan89]. formulae [IA85, IP89]. Fortran [Ada89]. forward [Rav87]. four [EE89]. Fourier [Car88b, KL89, Lon87c]. fractal [Mil88b]. Fredholm [NN87]. frontier [DG87a]. Fujitsu [LMM86]. full [Hei88]. function [Jer88, LE85, Mil88a, Uen89]. functional [DT87]. functions [Boy89, Ge89b, Ge89c, Hol89, IA85, Ioa88, KKS87]. fusion [Lar86].

Galerkin [Gol88b]. Game [NS89].

Game-dynamical [NS89]. Gauss [KL89]. general [Fah88, MM87, MS87b, Szi89, WKTM88]. Generalized [DT87, Gol88a]. generation [CSR88, Hof88, RS86]. generators [And87]. Geometric [Zak89, AB87]. geometrical [Eis87]. geometry [Mil88b]. Georgia [Tur88]. give [Yse86]. given [DG87b, Mil87]. Global [KB89, Sos88, Ge89c, Par89]. Gordon [BYPRV86]. governed [PM89]. gradient [Kap88, MKJ87, Yse86]. gravitational [BCZ88]. grazing [Fis85]. grid [CSR88, HMOT86, Hof88, RS86, Shy87]. grids [And87]. group [SR88]. groups [Pig87].

habitats [Fah88]. Hammerstein

[HL88b, HL88a]. hard [GS86]. Harmonic [Ioa88]. helium [Gre87b]. Helmholtz [JPS86]. Hierarchical [Yse86]. hierarchy [Har87, SV85]. high [DS86, Hol89]. higher [Mei89, Whi89, bS89]. higher-order [Whi89]. Hitachi [LMM86]. Hodgkin [Lam86]. homogeneous [MS87b]. homotopy [Mor86a, MS87a, MS87b]. Hopf [MR89]. horizon [Rav87]. Huxley [Lam86]. hybrid [Uen89]. hydrocode [KH87]. hydrodynamical [HK89]. hydrodynamics [BCZ88]. hydrologic [BW88]. hyperbolic [Car88a]. hypergames [WHF89]. hyperspheres [Lan89].

ideas [KP85]. identification [MMP89b]. II [Udw87b]. ill [Kee89]. ill-defined [Kee89]. image [Cha89]. imbedding [WKTM88]. impact [Fra89]. implicit [Bar89]. impulses [KB89]. impulsive [LL89, LP89, RSS89]. incidence [TT89]. incomplete [Axe86, Kap88]. incompressible [BS86]. Increasing [Lon87b]. indefinite [Man86]. index [Ano86a, Ano89g]. inequalities [AW87]. inference [MP86]. infinite [GW89]. infinity [Mor86b]. information [GFM88, Gil89]. initial [Kad86]. inputs [Pat87]. integer [GH89]. integrability [Mur89b]. integral [Bee87, Gol88b, Gol89, IP89, Kad86, KKS87, NN87, Sug87, TTB88]. integral-equation [Sug87]. Integrals [BE85, EE89, GW89]. integration [Beh89, bS89]. integrodifferential [Bee87, ER88, HL88b, HL88a, RSS89]. intensity [Jam86]. Interactive [DG87b]. interpolation [Beh89, Sai88]. Interspecific [Les89]. Interval [DTW88, MW88, Ge89a, Jam85, LS87, Wol89]. Introduction [DP86, Hic86, Jam88, CS89]. invariant [WKTM88]. invariant-imbedding [WKTM88]. invasions [CV89]. inventory [Zha89]. inverse [CAP88]. investigation [Ado87]. involving [RMR86, ZA88]. irrational [EE89]. issue [CS89]. iterated [Ber89]. iterative [Pan88, Sca89].

kernels [KKS87, NN87]. kind [NN87].

L [CG88]. L-shaped [CG88]. Lagrange [Got85, IA89]. Lagrangian [MM85]. Lakshmikantham [Zq87]. land [Tur88]. lands [BDT86]. landscape [CJ88, Jam88]. landscapes [Mil88b]. large [BW88, Ge89b, Zel87]. large-scale [BW88]. lattice [Mur89a]. least [Ber89, IA89, Sha89]. least-squares [Sha89]. lecturers [LH86]. Leela [Zq87]. Legendre [Lon87c]. less [Zem86]. less-developed [Zem86]. Levin [GW89]. Linear [DS86, Ada89, Ge89a, HL86, LR85, LV87, RK87, Sha89, Smi86, Sug87, WKTM88, Whi89]. lines [Ioa88, Mik87]. Lipschitz [ER88]. list [LH86]. locally [KM86]. log [Smi86]. log-linear [Smi86]. logarithmic [Boy89]. Losses [DDS85]. Low [BR88]. lower [LS87]. machines [Low86]. machining [MCD89]. macro [Zha89]. MACSYMA [RK87]. magnetic [VJ88]. maintenance [Ree87]. making [MCD89]. management [Ass87]. many [Kun86]. MAPPACC [Mat86]. mappings [Boy89]. market [Sos88]. marketing [Zem86]. Markov [Ben89, PM89]. Mathematical [Ger85, BCZ88, CS89, GD87]. Mathematics [Ano89a, Les89]. matrices [DLR85]. matrix [Axe86, Har87, KPT87, Sas85, Sha89]. max [MS88]. maximal [Pig87]. maximal-order [Pig87]. maximum [Rav87]. Maxwell [Cha88]. Mean [Sha87]. means [Mat86]. measurement [JB89a]. measurements [Jam85]. measures [LL89]. Measuring [Mil88b]. mechanics [Gre87a]. media [Cha88]. medication [Ass87]. medium [Sha87]. metabolism [Cas88]. metabolism-repair [Cas88]. Method [AG85, LV87, Ado87, BM89, BS86, BCZ88, CG87, Cha88, DR88, Fei85, Gol88b, Gre86, HMOT86, HL86, KR87a, KKS87, KM87, KL89, LS87, Mik87, NN87, Oma86, SR88, Shy87, bS89]. Methodology [Fis85]. methods [Aga86, Bee87, DTW88, Gol89, Hei88, Ier86, KB85, Kad86, Kap88, KW86, MM87, Man88, MMP89b, Pan88, SS86, Wol89, Yse86, ZH86]. MG [RS86]. MG-FAS [RS86]. min [MS88]. min-max [MS88]. minimizing [Ge89b]. Mixed [MS88, LV87]. mixture [BS86]. mixtures [BE85]. Model [EB86, BDT86, BCZ88, BCZ89, Con89, Fah88, Fra89, Jam85, KP86, Mos88b, Shi89, Sos88, SS87, TW89, Tur88, VJ86, Zha89]. Modeling [SV85, FMR86, Gre87b, Hen86, Hog88, Pal86]. models [Fra89, Kee89, MP86, NH86, Whi89]. modified [Bar89, WC88]. molecular [Gre87a, Kun86]. MOLP [DG87b]. moments [Lon87a]. monitoring [Jam86, Mos88a]. monotone [DR88, Oma86, Pan88]. monotony [LV87].

MOSI [Bar89]. Motion [VJ88, BS86]. MP/24 [LMM86]. mudpack [Ada89]. multi [LR85]. multi-time-scale [LR85]. multiattribute [BDT86]. multichannel [Ais88]. multicriteria [MCD89]. Multicriterion [Mat86]. Multigrid [GL86, Kam88, MM87, Man86, SS86, Axe86, Bar88, Bra86, Den86, Den88, Gre86, Hei88, HT88, Hol86, JPS86, KP85, KW86, KM87, KM86, KL89, Man88, Mel86, Mer86, Par85, Par87, TT86, TT88, Yse86, ZH86, Ada89]. Multigrid-sinc [SS86]. Multilevel [Mit86, Zem86]. multinormal [BE85]. Multiobjective [DDS85, HDP85, SS87]. multiparameter [pLR87b]. multiple [Cha88, CHMH86, DS87, JB89b]. multiplicative [BM89]. multiplicity [Shi89]. multipoint [AG85]. multiprocessors [Gre86]. multitasking [SM88]. Multivariate [JB89a, AB87, JB89b]. mutual [GFM88, Gil89].

n [HL86]. n-space [HL86]. Navier [Shy87, SJK86, ZH86]. neighborhoods [Mil87]. nested [KPT87, MP86]. net [Pat89]. network [CS85]. networks [Zel87, Zem86]. neutral [MB88]. Newton [DTW88]. no [Mur89b]. noise [TW89]. non [MP86]. non-nested [MP86]. Nonlinear [Cha89, Ado87, Ado88, Cas87, DT87, ER88, GH89, GW89, Hol89, Ier86, IP89, JPS86, LV87, Mit86, SW89]. Nonnormal [GD87]. nonrectangular [KB85]. nonstiff [bS89]. nonsymmetric [Man86]. norm [GP88]. Note [Gre87a, Gil89, Gol88a, HT88, Ott86, Par85, TT88, Zq87]. number [MM85]. Numerical [BYPRV86, CS85, Kad86, KR87b, KR87c, MMP89a, BS86, BCZ88, BCZ89, HK89, JPS86, KR87a, KR89, MM85, NN87, UA88, VJ88].

objectives [DG87b]. obtain [Mei89]. octonions [Car88a]. ODEs [bS89]. oil [VJ86]. one [DG87b, Man86, DG87b]. onset [DS87]. operations [Ger85, KPT87, MCD89]. operator [Bar89, SR88]. operators [GP88, TT86, Wol89]. Opportunity [DDS85]. Optimal [BdA89, Ge89a, Mos88b, Mos88a, Ree87, Fei85, Fis85, HH85, JB89b]. Optimization [FGLF88, Cas87, Ge89c, Jer88, JB89a, MS88, Par89, Pet88, SS87]. optimum [AB87]. order [CG87, HL88a, LS87, Nie85, Oma86, Pig87, Sas85, Whi89, Wol89, bS89]. order-interval [Wol89]. orders [DG87b]. organisms [Mei89]. oriented [Gre87a]. Oscillation [MB88]. oscillations [Zha89]. other [Ioa88].

Painlevé [VJ88]. paired [HQU88]. pairwise [Mat86]. para [Gre87b]. parabolic [MB88]. paradigm [Hog88]. Parallel [CJ88, CHMH86, MHL86, Par89, Pop86, SJK86, Axe86, Ge89b, Ge89c, GS86, Jam88, KM86, Nic86, TT88]. parallelism [Low86, Ott86]. parallelizable [HL86]. Parallelizing [OM86]. Parameter [CSR88, MMP89b, Bee87, Mit86, MS89]. parameters [Kee89, PM88, Pat89, Zia89]. parametrized [NH86]. part [LV87, RK87]. partial [Ada89, Bee87, DG87b, Mik87]. partially [BM89]. patchy [Fah88]. pattern [Mei89]. Patterns [CV89, MKJ87]. penalty [Fei85]. pendulum [Pet88]. performance [DS86, JB89a, LMM86]. Period [Pet88]. Periodic [Gae87, HL88b, HL88a, Den88]. Periodicity [TT89, BJ85]. Perron [Har87]. personal [Hic86]. perspective [Hic86]. perturbation [KR87a, KR87b, KR89, Kam88]. perturbed [DTW88, Gol88b, GP88, KR87c, LR85]. phenomena [HC87a]. physical [BJ85]. Physics [Con89, Ado87, Lar86]. physiology [AG89]. piecewise [AW87]. Piedmont [Tur88]. pipe [Sca89]. plan [FGLF88]. plane [Ioa88]. planning [CAP88, MCD89]. plant [Ger85]. plate [Whi89]. point [KR87c, UA88, WKTM88]. Pointwise [AG89, Eis87]. Poisson [And87, Bar88, Hol86]. polar [Bar88]. policies [Fis85, Mos88b]. polygon [Sai88]. polynomial [AP87, MW88, Mor86a, Mor86b, MS87a, MS87b, MS89]. population [Les89]. populations [Fah88, PM89]. portable [Ada89]. positive [Har87, MMP89a, Smi86]. Power [cC87, BdA89]. Power-series [cC87]. preconditioner [JPS86]. Preface [LH86, McC86]. preferences [Mat86]. presence [JB89b]. pressure [MKJ87]. pressure-gradient [MKJ87]. preventive [Ree87]. principle [IA89, Rav87]. priori [Jam85]. prisoner [NS89]. probabilities [Ben89, Smi86]. problem [BCZ89, CAP88, DG87b, HL88b, HL88a, Kun86, KH87, MS88, RMR86, SW89]. Problems [DDS85, AG85, Aga86, AB87, AG89, BM89, BJ85, Cas87, CK86, CHMH86, Den88, DTW88, Fei85, GS86, HK89, Ier86, KR87a, KR87b, KR87c, KR89, Kam88, KW86, MM87, Man86, Man88, Mit86, Oma86, Pan88, Rug86, TT88, UA88, WKTM88]. procedure [CG88]. procedures [MW88, TT86]. process [Har87, SV85]. processes [PM89]. processing [CJ88, CHMH86, Jam88]. processors [GS86]. production [SS87]. products [BE85]. Program [VA88]. programming [AB87, GH89, Ger85, GD87]. programs [Nic86, Ott86]. progress [KH87]. projection [Gol89]. projective [HL86]. propagation [Ber89]. properties [pLR87a]. protection [Ree87, Zam88]. prototype [Zam88]. Purcell [HL86].

quadrature [Ioa88]. quarry [Ger85]. quasilinear [Bee87, Pan88]. quasilinearization [WC88]. quaternions [Car88a]. queueing [Sos88]. Quintic [AP87]. quite [IP89]. radiation [CAP88]. radiation-therapy [CAP88]. radiative [KKS87]. radiotherapy [FGLF88]. Randles [LE85]. random [Jer88, Mos88a]. rate [Lon87b]. rational [Ge89c]. reaction [Mei89]. reaction-diffusion [Mei89]. reactor [Shi89]. real [DLR85]. reciprocal [Har87]. Reclamation [BDT86]. recovery [VJ86]. red [KL89]. red-black [KL89]. reflector [Uen89]. region [RMR86]. regionalized [CS85]. regions [CG88, Fis85, KB85]. Regression [Pat87, Pat89, Fra89]. Regular [NN87]. regulator [Sug87]. related [BCZ89]. relations [Mur89b]. relationship [Ott86]. relative [Ahm87, DS87]. Remark [Shi89]. Remarks [Par87]. repair [Cas88]. replacement [Ree87]. representation [Sug87]. representations [BJ85, Smi86]. reproducible [Mei89]. reservoir [VJ86]. residue [EE89]. resource [Les89]. respects [MS87b]. Response [Udw87a, Udw87b, VB89]. restoration [Cha89]. restrictions [MS88]. result [Gol88a]. results [HQU88, Rav87]. resumming [Lan89]. revenue [Ree87]. revenue-earning [Ree87]. reversible [LE85]. reversion [cC87]. revisited [Ahm87]. Riccati [BdA89, Sas85, Sas88]. root [Har87]. roots [DLR85, IP89, Mil87]. rotating [HH85]. rules [Ioa88]. run [Zha89].

S810 [LMM86]. S810/20 [LMM86]. sample [GFM88, Gil89]. Sampling [Jam86, AB87]. satellite [HH85]. scalar [Oma86]. scale [BW88, Ge89b, LR85]. scales [CHMH86]. scatterers [Sha87]. scattering [Uen89]. scheme [Bar89, DG87b, Ge89b, Pan88, Szi89, VJ88]. schemes [And87]. Schwarz [CG88]. scientific [Nic86]. search [Jer88, Par89]. Searching [Hol89]. second [HL88a, NN87, Oma86, Sas85]. sedenions [Car88a]. Seidel [KL89]. selection [EB86, Fra89]. semi [PM89]. semi-Markov [PM89]. Semiclassical [Gre87b]. semiderivatives [LCE85]. semiintegrals [LCE85]. Sensitivity [EBR85, PM88]. separable [Ge89c]. separable-factorable [Ge89c]. sequences [HC87b]. series [BJ85, Car88b, cC87, FMR86, GW89, Lon87a, Lon87b, Lon87c, Pal86]. series-solution [BJ85]. service [Pat87, PM88]. set [Ahm87, Nie85, Pig87]. sets [KB89, Sos88]. Sevcik [LE85]. several [Sca89]. shaped [CG88]. ship [Zam88]. shock [HK89, Zak89]. shocks [Arn86]. Short [Zha89, TT88]. Short-run [Zha89]. shortage [Sos88]. Sidi [GW89]. simple [IA85, Pet88]. Simulating [Arn86]. simulation [BW88, Tur88, Zam88]. simultaneous [MW88]. sinc [SS86]. sine [BYPRV86]. sine-Gordon [BYPRV86]. single [RS86]. single-equation [RS86]. singlets [Gre87b]. singular [AG89, Boy89, CS88, Den88, Gol88b, Gol89, Kad86, KR87a, KR87b, KR89, Kam88, LV87]. singularities [Boy89]. Singularity [Cas87]. singularly [GP88, KR87c, LR85]. size [GFM88, Gil89]. smoothers [Axe86, MM87]. smoothing [KL89, Man86, TT86]. software [Ada89]. solar [TT89]. Solution [WHF89, Ada89, AP87, AG89, BYPRV86, Ber89, BJ85, CAP88, Cha88, Ier86, JPS86, Kad86, KR87a, KR87b, MMP89a, Mik87, NN87, Nie85, Sca89, Sha87, UA88, WKTM88]. Solutions [Lam86, AR87, Ado88, HK89, LS87, MB88, Mor86b, MS87a, Mur89a, Oma86, RSS89, Sha89]. solver [Bar88, Bar89, HL86, Hol86, Mer86, RS86]. Solving [bS89, Fei85, KB85, Kun86, Mor86a, MS87b, SR88]. Some [HQU88, KH87, SM88, Bee87, GS86, KP85, MW88]. space [Ahm87, HL86]. spaces [Vál88]. span [Hof88]. sparse [HL86]. spatial [DS87, Tur88]. Special [CS89]. specification [KP86]. spectra [Gre87b]. Spectral [Ier86, ZH86]. speed [Yse86]. spherical [RMR86]. SPICE [OM86]. spline [Beh89]. spread [Gle88, Wat89]. squares [Ber89, Sha89]. Stability [Ahm87, LL89, ER88, KB89, LR85, LP89, Mik87, Sos88]. stabilizing [HH85]. Stable [CV89]. stages [Ass88]. standard [TT88]. State [BL87, PM89]. states [Gre87b]. Statistical [MP86, PM89]. Stefan [RMR86]. step [Man86]. Stochastic [Uen89, CK86, GD87, JPS86, PM88, Pat89]. stocking [Fis85]. Stokes [Shy87, SJK86, ZH86]. Strategies [BW88, DG87a, KP86, Mos88a, CV89]. strategy [Jer88]. Strict [Vál88]. stripmined [BDT86]. strong [Arn86]. strongly [Bar89]. Structural [FMR86, Pal86, Kee89]. structure [Nie85, Zak89]. structures [MS87b]. Study [Mur89a, Ier86, KP85, Man88, TW89]. subgrid [RS86]. Subject [Ano89g]. subjected [Mos88a]. subvortices [GL86]. successive [CK86]. sufficient [Shi89]. sum [MMP89a]. summation [Lon87a, Lon87c]. supercomputer [KN86]. supercomputers [LMM86]. superconvergence [Gol88a]. supergrid [RS86]. survey [KR89]. symmetric [Bra86, MM87, Man88, Sca89]. synthesis [HH85]. system [Ais88, Ass87, Ass88, MR89, Pat87, PM88, SM88, Uen89, Zam88]. systems [BdA89, BL87, Cas88, CG87, Den86, ER88, Jam86, Kee89, KM86, KB89, Kun86, LR85, LL89, LP89, pLR87b, Mei89, Mor86a, Mor86b, MS87a, MS87b, RSS89, Udw87a, Udw87b, VB89].

Tailoring [Mei89]. Tam [Lan89]. tapered [Hof88]. Taylor [BR88]. techniques [Mit86, MCD89, Sca89]. terminal [BM89]. termination [VA88]. terms [LL89, Whi89]. test [Jam85]. Testing [KP86, NH86]. theorem [Got85, Sas88, Zq87]. theorems [Par87, Smi86]. theory [Bra86, Cas87, Cas88, EE89, MM87, Mur89a, Mur89b, Pet88, Shi89, Szi89, VA88]. therapy [CAP88]. three [Arn86, CG87, Hol86, Kap88, KW86, KM87, LMM86, ZA88]. three-dimensional [Hol86, KM87, ZA88]. time [FMR86, HH85, LR85, Pal86, PM89, ZH86]. time-dependent [ZH86]. timesharing [OM86]. Topological [Zel87]. transfer [KKS87]. transform [VA88, WH85]. transformation [GD87, Mor86b]. transformations [GW89]. transition [Ben89]. transonic [Mel86]. transonic-flow [Mel86]. transport [Zia89]. treatment [CAP88, FGLF88, KR87c, TT86]. Truncated [Kap88]. turbine [WC88]. turbulence [DS87]. Two [KL89, Smi86, Ass88, BS86, DG87b, KR87c, Kap88, LL89, Sca89, UA88, WKTM88, Zia89]. Two-color [KL89]. two-dimensional [Zia89]. two-pipe [Sca89]. two-point [WKTM88]. type [HL88b, HL88a, MB88, RMR86, Yse86].

uncertain [BL87, Udw87a, Udw87b]. Uncertainty [DDS85]. unification [cC87]. unified [Con89]. unit [Bar88]. unweighted [Sha89]. upper [LS87]. Use [Mur89b, JB89a, Tur88]. using [Bar88, EE89, MS87a, NH86, Pat89, Pet88, RK87, VA88]. usual [MM87].

validity [Jam85, Mik87]. value [AG85, Aga86, CS88, HC87a, HL88b, HL88a, Ier86, Kad86, KR87c, pLR87a, pLR87b, Mit86, Oma86, Pan88, SW89, UA88, WKTM88]. values [HC87b, Hol89]. variable [Mos88a]. variables [CS85, DG87b]. variation [Bee87, DT87]. variation-of-constants [DT87]. variational [CSR88, MM87, Man86]. various [MKJ87]. vector [Axe86, CJ88, Jam88, NH86, Vál88]. Vectorizable [Mel86]. vectorization [HT88]. vectorized [Hol86]. velocity [MKJ87]. versions [MW88]. very [Arn86]. via [CS88, Kad86, KR87b]. viscous [BS86]. visual [Hof89]. Volterra [ER88]. Volume [Ano89a]. Volumes [Lan89, Ano86a, Ano89g]. VP [LMM86]. VP-200 [LMM86].

waiting [Ass88]. wave [Sha87, Soc88]. waves [Zak89]. weak [Zak89]. Weighted [Sha89, Pan88]. whole [Ahm87]. wine [SS87]. within [Pig87]. workshop [Jam88].

X [LMM86]. X-MP [LMM86]. X-MP/24 [LMM86].

Zardecki [Lan89]. zeros [IA85, Mil88a, MW88, Wol89].

References

Ahmed:1987:AGP

[AB87] Javed Ahmed and Charles D. Bonham. Application of geometric programming to optimum allocation problems in multivariate doubling sampling. Applied Mathematics and Computation, 21(2):157–169, February 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                                   tronic).  URL http://www.
                                                   sciencedirect.com/science/
                                                   article/pii/0096300387900245.

Adams:1989:MMP

[Ada89] John C. Adams. mudpack: Multigrid portable Fortran software for the efficient solution of linear elliptic partial differential equations. Applied Mathematics and Computation, 34(2 (part 2)): 113–146, November 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                                   tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0096300389900106.

Adomian:1987:IAD

[Ado87] G. Adomian. An investigation of the asymptotic decomposition method for nonlinear equations in physics. Applied Mathematics and Computation, 24 (1):1–17, October 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900270.

Adomian:1988:ASN

[Ado88] G. Adomian. Analytic solutions for nonlinear equations. Applied Mathematics and Computation, 26(1):77–88, April 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900872.

Agarwal:1985:MCM

[AG85] Ravi P. Agarwal and Ramesh C.

Gupta. Method of chasing for multipoint boundary value problems. Applied Mathematics and Computation, 17(1):37–43, July 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300385900177.

Asaithambi:1989:PSB

[AG89] N. S. Asaithambi and J. B. Garner. Pointwise solution bounds for a class of singular diffusion problems in physiology. Applied Mathematics and Computation, 30(3):215–222, April 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900532.

Agarwal:1986:CMD

[Aga86] Ravi P. Agarwal. Computational methods for discrete boundary value problems. Applied Mathematics and Computation, 18 (1):15–41, January 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386900263.

Ahmad:1987:SRS

[Ahm87] Khader H. Ahmad. Stability relative to a set and to the whole space revisited. Applied Mathematics and Computation, 24(2): 91–99, November 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900622.

Aisbett:1988:CAM

[Ais88] Janet Aisbett. Cyclic audit of a multichannel system. Applied Mathematics and Computation, 25(1):11–28, January 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900616.

Anderson:1987:ESP

[And87] Dale A. Anderson. Equidistribution schemes, Poisson generators, and adaptive grids. Applied Mathematics and Computation, 24(3):211–227, December 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900853.

Anonymous:1985:EB

[Ano85] Anonymous. Editorial board.

Applied Mathematics and Computation, 17(1):ii–iii, July 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300385900153.

Anonymous:1986:AIV

[Ano86a] Anonymous. Author index to volumes 18–20, 1986. Applied Mathematics and Computation, 20(3– 4):401–404, November 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386900135.

Anonymous:1986:EBa

[Ano86b] Anonymous. Editorial board.

Applied Mathematics and Computation, 18(1):ii–iii, January 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630038690024X.

Anonymous:1986:EBb

[Ano86c] Anonymous. Editorial Board. Applied Mathematics and Computation, 19(1–4):ii–iii, July 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300386900925.

Anonymous:1986:EBc

[Ano86d] Anonymous. Editorial Board. Applied Mathematics and Computation, 20(1–2):ii–iii, September 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/science/
                                        article/pii/0096300386901153.

Anonymous:1987:EBa

[Ano87a] Anonymous. Editorial Board. Applied Mathematics and Computation, 21(1):ii–iii, January 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300387900051.

Anonymous:1987:EBb

[Ano87b] Anonymous. Editorial Board. Applied Mathematics and Computation, 22(1):ii–iii, April 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900464.

Anonymous:1987:EBc

[Ano87c] Anonymous. Editorial Board. Applied Mathematics and Computation, 23(1):ii–iii, July 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900518.

Anonymous:1987:EBd

[Ano87d] Anonymous. Editorial Board.

Applied Mathematics and Computation, 24(1):ii–iii, October 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300387900269.

Anonymous:1988:EBa

[Ano88a] Anonymous. Editorial Board.

Applied Mathematics and Computation, 26(1):ii–iii, April 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900835.

Anonymous:1988:EBb

[Ano88b] Anonymous. Editorial board.

Applied Mathematics and Computation, 27(1):ii–iii, July 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900938.

Anonymous:1988:EBc

[Ano88c] Anonymous. Editorial Board.

Applied Mathematics and Computation, 28(1):ii–iii, October 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300388901294.

Anonymous:1989:AMC

[Ano89a] Anonymous. Applied Mathematics and Computation: Volume 34. Applied Mathematics and Computation, 34(3):i–ii, December 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300389901057.

Anonymous:1989:EBa

[Ano89b] Anonymous. Editorial Board. Applied Mathematics and Computation, 29(1):ii–iii, January 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300389900350.

Anonymous:1989:EBb

[Ano89c] Anonymous. Editorial Board. Applied Mathematics and Computation, 30(1):ii–iii, March 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300389900611.

Anonymous:1989:EBc

[Ano89d] Anonymous. Editorial Board. Applied Mathematics and Computation, 32(1):ii–iii, July 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0096300389900428.

Anonymous:1989:EBd

[Ano89e] Anonymous. Editorial Board.

Applied Mathematics and Computation, 33(1):ii–iii, September 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300389900568.

Anonymous:1989:EBe

[Ano89f] Anonymous. Editorial Board.

Applied Mathematics and Computation, 34(1):ii–iii, November 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300389900015.

Anonymous:1989:SIV

[Ano89g] Anonymous. Subject index to volumes 29–34, 1989. Applied Mathematics and Computation, 34(3):iii–vi, December 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389901069.

Anastasselou:1987:ASP

[AP87] E. G. Anastasselou and D. E.

Panayotounakos. Analytical solution of polynomial equations with an application to the Quintic equation. Applied Mathematics and Computation, 21(2): 145–155, February 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900233.

Adomian:1987:ESD

[AR87] G. Adomian and R. Rach. Explicit solutions for differential equations. Applied Mathematics and Computation, 23 (1):49–59, July 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900555.

Arnold:1986:SVS

[Arn86] Clifford N. Arnold. Simulating very strong shocks in three dimensions. Applied Mathematics and Computation, 20(1– 2):53, September 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300386901244.

Assimakopoulos:1987:MMS

[Ass87] Nikitas Assimakopoulos. A medication management system. Applied Mathematics and Computation, 21(1):73–91, January 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900105.

Assimakopoulos:1988:WST

[Ass88] Nikitas Assimakopoulos. A waiting system with two stages. Applied Mathematics and Computation, 27(2):173–190, August 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300388900288.

Aftabizadeh:1987:DID

[AW87] A. R. Aftabizadeh and Joseph Wiener. Differential inequalities for delay differential equations with piecewise constant argument. Applied Mathematics and Computation, 24(3): 183–194, December 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038790083X.

Axelsson:1986:AIM

[Axe86] O. Axelsson. Analysis of incomplete matrix factorizations as multigrid smoothers for vector and parallel computers. Applied Mathematics and Computation, 19(1–4):3–22, July 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386900949.

Barros:1988:PEU

[Bar88] Saulo R. M. Barros. The Poisson equation on the unit disk: a multigrid solver using polar coordinates. Applied Mathematics and Computation, 25(2): 123–135, January 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388901105.

Barthelson:1989:MOS

[Bar89] Stephen H. Barthelson. The modified operator strongly implicit scheme, MOSI, as a fast elliptic solver. Applied Mathematics and Computation, 33(3): 185–196, October 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300389900507.

Bulgarelli:1988:NMM

[BCZ88] U. Bulgarelli, M. M. Cerimele, and A. Zaretti. A numerical method for a mathematical model in gravitational hydrodynamics. Applied Mathematics and Computation, 27(3): 251–264, September 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388900069.

Bulgarelli:1989:NMR

[BCZ89] Ulderico Bulgarelli, Maria Mercede Cerimele, and Anna Zaretti. A numerical model related to an astrophysical problem. Applied Mathematics and Computation, 29(1):17–38, January 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0096300389900374.

Ballesteros:1989:OCP

[BdA89] F. Ballesteros and F. de Arriaga.

Optimal control of power systems and the Riccati equation. Applied Mathematics and Computation, 33(1):43–52, September 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300389900581.

Brinck:1986:RSL

[BDT86] Fritz H. Brinck, Lucien Duckstein, and John Thames. Reclamation of stripmined lands: a multiattribute decision model. Applied Mathematics and Computation, 18(2):167–183, February 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/
article/pii/0096300386900238.

Bonney:1985:IPM

[BE85] George E. Bonney and Robert C.

Elston. Integrals of products of multinormal mixtures. Applied Mathematics and Computation, 16(2):93–104, February 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300385900013.

Beesack:1987:SVP

[Bee87] Paul R. Beesack. On some variation of parameter methods for integrodifferential, integral, and quasilinear partial integrodifferential equations. Applied Mathematics and Computation, 22 (2–3):189–215, May 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900427.

Behforooz:1989:ECC

[Beh89] G. Behforooz. End conditions for cubic spline interpolation derived from integration. Applied Mathematics and Computation, 29(3): 231–244, February 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900155.

Benaroya:1989:MCT

[Ben89] Haym Benaroya. Markov chain transition probabilities and experimental data. Applied Mathematics and Computation, 29(2): 107–121, January 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900982.

Berman:1989:EPE

[Ber89] Sorrell Berman. An example of propagation of error associated with an iterated least squares solution. Applied Mathematics and Computation, 34(3): 173–193, December 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389901033.

Byrd:1985:DPP

[BJ85] J. William Byrd and Robert R.

Jones. The determination of periodicity in physical problems from series-solution representations. Applied Mathematics and Computation, 16(3):203–211, April 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300385900281.

Breinl:1987:SFU

[BL87] W. Breinl and G. Leitmann.

State feedback for uncertain dynamical systems. Applied Mathematics and Computation, 22 (1):65–87, April 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900506.

Bainov:1989:MPM

[BM89] D. D. Bainov and S. D. Milusheva. The method of partially multiplicative averaging in the problems of terminal control. Applied Mathematics and Computation, 33(1):69–74, September 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630038990060X.

Boyd:1989:ACC

[Boy89] John P. Boyd. The asymptotic Chebyshev coefficients for functions with logarithmic endpoint singularities: mappings and singular basis functions. Applied Mathematics and Computation, 29(1):49–67, January 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300389900398.

Bao:1988:LCD

[BR88] Paul G. Bao and Jon G. Rokne. Low complexity k-dimensional Taylor forms. Applied Mathematics and Computation, 27(3): 265–280, September 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388900070.

Brandt:1986:AMT

[Bra86] Achi Brandt. Algebraic multigrid theory: the symmetric case. Applied Mathematics and Computation, 19(1–4):23–56, July 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300386900950.

Bulgarelli:1986:NMM

[BS86] U. Bulgarelli and R. M. Spitaleri. A numerical method for the motion of a mixture of two viscous incompressible fluids. Applied Mathematics and Computation, 18(3):223–237, April 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386900160.

binSuleiman:1989:SNH

[bS89] Mohamed bin Suleiman. Solving nonstiff higher order ODEs directly by the direct integration method. Applied Mathematics and Computation, 33(3): 197–219, October 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900519.

Band:1988:SLS

[BW88] Lawrence E. Band and Eric F.

Wood. Strategies for large-scale, distributed hydrologic simulation. Applied Mathematics and Computation, 27(1):23–37, July 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300388900963.

Ben-Yu:1986:NSS

[BYPRV86] Guo Ben-Yu, Pedro J. Pascual, María J. Rodriguez, and Luis Vázquez. Numerical solution of the sine-Gordon equation. Applied Mathematics and Computation, 18(1):1–14, January 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300386900251.

Censor:1988:CSI

[CAP88] Yair Censor, Martin D. Altschuler, and William D. Powlis. A computational solution of the inverse problem in radiationtherapy treatment planning. Applied Mathematics and Computation, 25(1):57–87, January 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388900641.

Carmody:1988:CHQ

[Car88a] Kevin Carmody. Circular and hyperbolic quaternions, octonions, and sedenions. Applied Mathematics and Computation, 28 (1):47–72, October 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388901336.

Carmody:1988:FFS

[Car88b] Kevin Carmody. Finite Fourier series. Applied Mathematics and Computation, 28(2): 89–96, November 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388900884.

Casti:1987:STN

[Cas87] J. Casti. Singularity theory for nonlinear optimization problems.

Applied Mathematics and Computation, 23(2):137–161, August 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630038790035X.

Casti:1988:TMR

[Cas88] John L. Casti. The theory of metabolism-repair systems. Applied Mathematics and Computation, 28(2):113–154, November 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300388900902.

Chang:1987:PSU

[cC87] Feng cheng Chang. Power-series unification and reversion. Applied Mathematics and Computation, 23(1):7–23, July 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900531.

Chang:1987:FEM

[CG87] Ching-Lung Chang and Max D.

Gunzburger. A finite element method for first order elliptic systems in three dimensions. Applied Mathematics and Computation, 23(2):171–184, August 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900373.

Cai:1988:EES

[CG88] Zhiqiang Cai and Benju Guo. Error estimates for a Schwarz alternating procedure on Lshaped regions. Applied Mathematics and Computation, 28 (1):39–46, October 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388901324.

Chang:1988:FEM

[Cha88] Ching-Lung Chang. Finite element method for the solution of Maxwell’s equations in multiple media. Applied Mathematics and Computation, 25 (2):89–99, January 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388901075.

Chang:1989:NFI

[Cha89] Mou-Hsiung Chang. Nonlinear filtering for image restoration. Applied Mathematics and Computation, 33(2):131– 147, September 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300389900878.

Chin:1986:PPM

[CHMH86] R. C. Y. Chin, G. W. Hedstrom, J. R. McGraw, and F. A. Howes. Parallel processing of multiple scales problems. Applied Mathematics and Computation, 20(1– 2):55, September 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386901256.

Casey:1988:PVP

[CJ88] Richard M. Casey and Donald A.

Jameson. Parallel and vector processing in landscape dynamics. Applied Mathematics and Computation, 27(1):3–22, July 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900951.

Chang:1986:SAA

[CK86] M. H. Chang and K. Krishna.

A successive approximation algorithm for stochastic control problems. Applied Mathematics and Computation, 18(2): 155–165, February 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386900226.

Conrad:1989:PBT

[Con89] Michael Conrad. Physics and biology: towards a unified model. Applied Mathematics and Computation, 32(2–3):75–102, August 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300389900891.

Carrera:1985:NCN

[CS85] Jesus Carrera and Ferenc Szidarovszky. Numerical comparison of network design algorithms for regionalized variables. Applied Mathematics and Computation, 16(3):189–202, April 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630038590027X.

Chow:1988:BSV

[CS88] Shui-Nee Chow and Yun-Qiu Shen. Bifurcations via singular value decompositions. Applied Mathematics and Computation, 28(3):231–245, December 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388901397.

Casti:1989:SIM

[CS89] John Casti and Karl Sigmund. Special issue on mathematical biology: Introduction. Applied Mathematics and Computation, 32(2–3):73–74, August 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630038990088X.

Castillo:1988:PEV

[CSR88] José Castillo, Stanly Steinberg, and Patrick J. Roache. Parameter estimation in variational grid generation. Applied Mathematics and Computation, 28(2): 155–177, November 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900914.

Cannings:1989:PIE

[CV89] C. Cannings and G. T. Vickers.

Patterns and invasions of Evolutionary Stable Strategies. Applied Mathematics and Computation, 32(2–3):227–253, August 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900957.

Davis:1985:OLD

[DDS85] D. Davis, L. Duckstein, and F. Szidarovszky. Opportunity losses due to uncertainty in multiobjective decision problems. Applied Mathematics and Computation, 17(1):69–91, July 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300385900190.

Dendy:1986:BBM

[Den86] J. E. Dendy, Jr. Black box multigrid for systems. Applied Mathematics and Computation, 19(1–4):57–74, July 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300386900962.

Dendy:1988:BBM

[Den88] J. E. Dendy, Jr. Black box multigrid for periodic and singular problems. Applied Mathematics and Computation, 25 (1):1–10, January 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388900604.

Deckro:1987:DSF

[DG87a] Richard F. Deckro and Thomas M. Gunther. Drilling strategies for frontier areas. Applied Mathematics and Computation, 23 (1):71–91, July 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900579.

Dror:1987:ISM

[DG87b] Moshe Dror and Saul I. Gass. Interactive scheme for a MOLP problem given two partial orders: One on variables and one on objectives. Applied Mathematics and Computation, 24(3): 195–209, December 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900841.

Denman:1985:CRR

[DLR85] E. D. Denman and J. Leyva-Ramos. Computation of roots of real and complex matrices. Applied Mathematics and Computation, 16(3):213–228, April 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300385900293.

Deistler:1986:I

[DP86] M. Deistler and B. M. Pötscher.

Introduction. Applied Mathematics and Computation, 20(3–4): 197–199, November 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386900056.

Devasahayam:1988:MME

[DR88] M. Paul Devasahayam and V. Subba Rao. A monotone method for the equation u_{xyz} = f. Applied Mathematics and Computation, 25(3): 175–186, February 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900707.

Dongarra:1986:LAH

[DS86] J. J. Dongarra and D. C.

Sorensen. Linear algebra on high performance computers. Applied Mathematics and Computation, 20(1–2):57–88, September 1986.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386901268.

Dendrinos:1987:OTD

[DS87] Dimitrios S. Dendrinos and Michael Sonis. The onset of turbulence in discrete relative multiple spatial dynamics. Applied Mathematics and Computation, 22(1):25–44, April 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900488.

Deo:1987:GVC

[DT87] Sadashiv G. Deo and Euclides F. Torres. Generalized variationof-constants formula for nonlinear functional differential equations. Applied Mathematics and Computation, 24(3):263– 274, December 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900889.

Dinkel:1988:INM

[DTW88] John J. Dinkel, Marietta Tretter, and Danny Wong. Interval Newton methods and perturbed problems. Applied Mathematics and Computation, 28(3): 211–222, December 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0096300388901373.

Engle:1986:MSF

[EB86] Robert F. Engle and Scott J.

Brown. Model selection for forecasting. Applied Mathematics and Computation, 20(3–4): 313–327, November 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386900093.

Eno:1985:SAE

[EBR85] Larry Eno, Johan G. B. Beumee, and Herschel Rabitz. Sensitivity analysis of experimental data. Applied Mathematics and Computation, 16(2): 153–163, February 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300385900050.

Evans:1989:EFI

[EE89] James D. Evans and Lori M.

Evans. Evaluation of four irrational cosine definite integrals using residue theory. Applied Mathematics and Computation, 33(3): 149–159, October 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900489.

Eiseman:1987:GCP

[Eis87] Peter R. Eiseman. The geometrical construction of pointwise distributions on curves. Applied Mathematics and Computation, 21(3):233–275, February 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630038790004X.

Elaydi:1988:LSN

[ER88] S. Elaydi and M. Rama Mohana Rao. Lipschitz stability for nonlinear Volterra integrodifferential systems. Applied Mathematics and Computation, 27(3): 191–199, September 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630038890001X.

Fahrig:1988:GMP

[Fah88] L. Fahrig. A general model of populations in patchy habitats. Applied Mathematics and Computation, 27(1):53–66, July 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388900987.

Feiring:1985:EPM

[Fei85] Bruce R. Feiring. An exact penalty method for solving optimal control problems. Applied Mathematics and Computation, 16(2):105–114, February 1985.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300385900025.

Fymat:1988:ORT

[FGLF88] A. L. Fymat, M. A. Greenfield, S. C. Lo, and D. Findley. Optimization of the radiotherapy treatment plan. Applied Mathematics and Computation, 27 (2):117–146, August 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900240.

Fisher:1985:DOS

[Fis85] I. H. Fisher. Derivation of optimal stocking policies for grazing in arid regions. I. Methodology. Applied Mathematics and Computation, 17(1):1–35, July 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300385900165.

Florens:1986:STS

[FMR86] J.-P. Florens, M. Mouchart, and J.-F. Richard. Structural time series modeling: a Bayesian approach. Applied Mathematics and Computation, 20(3–4): 365–400, November 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386900123.

Franses:1989:DBR

[Fra89] Philip Hans Franses. The distance between regression models and its impact on model selection. Applied Mathematics and Computation, 34(1): 1–16, November 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300389900027.

Gaertner:1987:PAC

[Gae87] Wulf Gaertner. Periodic and aperiodic consumer behavior. Applied Mathematics and Computation, 22(2–3):233–254, May 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900440.

Goicoechea:1987:NDE

[GD87] Ambrose Goicoechea and Lucien Duckstein. Nonnormal deterministic equivalents and a transformation in stochastic mathematical programming. Applied Mathematics and Computation, 21(1):51–72, January 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900099.

Ge:1989:OCL

[Ge89a] Renpu Ge. Optimal choice of linear interval extension. Applied Mathematics and Computation, 30(2):165–189, March 1989.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389901495.

Ge:1989:PCS

[Ge89b] Renpu Ge. A parallel computing scheme for minimizing a class of large scale functions. Applied Mathematics and Computation, 30(3):261–288, April 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900556.

Ge:1989:PGO

[Ge89c] Renpu Ge. A parallel global optimization algorithm for rational separable-factorable functions. Applied Mathematics and Computation, 32(1):61–72, July 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300389900477.

Gershon:1985:MPA

[Ger85] Mark E. Gershon. Mathematical programming applied to cementplant and quarry operations. Applied Mathematics and Computation, 16(2):165–176, February 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300385900062.

Gil:1988:CSS

[GFM88] M. A. Gil, M. J. Fernández, and I. Martínez. The choice of sample size in estimating mutual information. Applied Mathematics and Computation, 27(3): 201–216, September 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388900021.

Ge:1989:CAN

[GH89] Renpu Ge and Changbin Huang. A continuous approach to nonlinear integer programming. Applied Mathematics and Computation, 34(1):39–60, November 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300389900052.

Gil:1989:NCS

[Gil89] María Angeles Gil. A note on the choice of sample size in estimating the mutual information. Applied Mathematics and Computation, 30(2):125–132, March 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630038990146X.

Gustafson:1986:MCS

[GL86] Karl Gustafson and Robert Leben. Multigrid calculation of subvortices. Applied Mathematics and Computation, 19 (1–4):89–102, July 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386900986.

Gleissner:1988:SE

[Gle88] Winfried Gleißner. The spread of epidemics. Applied Mathematics and Computation, 27(2): 167–171, August 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900276.

Golberg:1988:NSR

[Gol88a] Michael A. Golberg. A note on a superconvergence result for the generalized airfoil equation. Applied Mathematics and Computation, 26(2):105–117, May 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900458.

Golberg:1988:PGM

[Gol88b] Michael A. Golberg. The perturbed Galerkin method for Cauchy singular integral equations with constant coefficients. Applied Mathematics and Computation, 26(1):1–33, April 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900847.

Golberg:1989:DPM

[Gol89] Michael A. Golberg. Discrete projection methods for Cauchy singular integral equations with constant coefficients. Applied Mathematics and Computation, 33(1):1–41, September 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630038990057X.

Gotusso:1985:ETL

[Got85] Laura Gotusso. On the energy theorem for the Lagrange equations in the discrete case. Applied Mathematics and Computation, 17(2):129–136, September 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300385900232.

Goldstein:1988:NES

[GP88] Charles I. Goldstein and Seymour V. Parter. On the norm equivalence of singularly perturbed elliptic difference operators. Applied Mathematics and Computation, 26(3): 179–200, May 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388901191.

Greenbaum:1986:MMM

[Gre86] A. Greenbaum. A multigrid method for multiprocessors. Applied Mathematics and Computation, 19(1–4):75–88, July 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386900974.

Greenspan:1987:NNC

[Gre87a] Donald Greenspan. Note on a new computer-oriented approach to molecular mechanics. Applied Mathematics and Computation, 21(1):45–50, January 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900087.

Greenspan:1987:SMS

[Gre87b] Donald Greenspan. Semiclassical modeling of the states and spectra of para helium singlets with an algebraic formula for excited states. Applied Mathematics and Computation, 24(3): 229–240, December 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900865.

Gilpin:1986:APP

[GS86] Michael J. Gilpin and Robert O.

Shelton. Applications of parallel processors to some hard discrete problems. Applied Mathematics and Computation, 20(1– 2):1–4, September 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300386901177.

Gray:1989:ELS

[GW89] Henry L. Gray and Suojin Wang. An extension of the Levin–Sidi class of nonlinear transformations for accelerating convergence of infinite integrals and series. Applied Mathematics and Computation, 33(2):75–87, September 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300389900842.

Harker:1987:DPR

[Har87] Patrick T. Harker. Derivatives of the Perron root of a positive reciprocal matrix: With application to the analytic hierarchy process. Applied Mathematics and Computation, 22 (2–3):217–232, May 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900439.

Hirtzel:1987:CEA

[HC87a] C. S. Hirtzel and K. Y. Chan. Computer experiments for the analysis of extreme-value phenomena. Applied Mathematics and Computation, 21(2): 125–144, February 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0096300387900221.

Hirtzel:1987:EVE

[HC87b] C. S. Hirtzel and K. Y. Chan.

Extreme values of exponentially autocorrelated sequences. Applied Mathematics and Computation, 22(1):1–24, April 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900476.

Hiessl:1985:MAC

[HDP85] H. Hiessl, I. Duckstein, and E. J.

Plate. Multiobjective Q-analysis with concordance and discordance concepts. Applied Mathematics and Computation, 17(2): 107–122, September 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300385900219.

Heinrichs:1988:CFM

[Hei88] Wilhelm Heinrichs. Collocation and full multigrid methods. Applied Mathematics and Computation, 26(1):35–43, April 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900859.

Hendry:1986:EMD

[Hen86]  David F. Hendry.  Empiri-

cal modeling in dynamic econometrics. Applied Mathematics and Computation, 20(3–4): 201–236, November 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386900068.

Hermes:1985:ESS

[HH85] Henry Hermes and David Hogenson. The explicit synthesis of stabilizing (time optimal) feedback controls for the attitude control of a rotating satellite. Applied Mathematics and Computation, 16(3):229–240, April 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038590030X.

Hicks:1986:ICC

[Hic86]  Darrell L. Hicks.  Introduc-

tion to computational continuum dynamics: a personal perspective. Applied Mathematics and Computation, 20(1–2): 5–19, September 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386901189.

Hicks:1989:CNS

[HK89] D. L. Hicks and K. L. Kuttler. The convergence of numerical solutions of hydrodynamical shock problems. Applied Mathematics and Computation, 30(2):99–123, March 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389901458.

Hill:1986:PMP

[HL86] Richard O. Hill, Jr. and Gerald D. Ludden. Purcell’s method, projective n-space, and a parallelizable, sparse linear equation solver. Applied Mathematics and Computation, 20(1–2): 89–94, September 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038690127X.

Hu:1988:PBVb

[HL88a] Shouchuan Hu and S. Leela.

Periodic boundary value problem for a second order integrodifferential equation of Hammerstein type. Applied Mathematics and Computation, 25(3): 219–226, February 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900732.

Hu:1988:PBVa

[HL88b] Shouchuan Hu and S. Leela. Periodic boundary value problem for integrodifferential equations of Hammerstein type. Applied Mathematics and Computation, 25(1):29–38, January 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300388900628.

Hart:1986:FAC

[HMOT86] L. Hart, S. McCormick, A. O’Gallagher, and J. Thomas. The fast adaptive composite-grid method (FAC): algorithms for advanced computers. Applied Mathematics and Computation, 19(1– 4):103–125, July 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300386900998.

Hoffman:1988:AGG

[Hof88] Gilbert H. Hoffman. Algebraic grid generation for an afterbody with finite-span, tapered fins. Applied Mathematics and Computation, 28(1):15–37, October 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300388901312.

Hoffman:1989:VCC

[Hof89] William C. Hoffman. The visual cortex is a contact bundle. Applied Mathematics and Computation, 32(2–3):137–167, August 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630038990091X.

Hogeweg:1988:CAP

[Hog88] P. Hogeweg. Cellular automata as a paradigm for ecological modeling. Applied Mathematics and Computation, 27(1):81–100, July 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300388901002.

Holter:1986:VMS

[Hol86] William H. Holter. A vectorized multigrid solver for the threedimensional Poisson equation. Applied Mathematics and Computation, 19(1–4):127–144, July 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300386901001.

Holland:1989:SNF

[Hol89]  John H. Holland.  Search-

ing nonlinear functions for high values. Applied Mathematics and Computation, 32(2–3): 255–274, August 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900969.

Hanumara:1988:SRB

[HQU88] R. Choudary Hanumara, Eugene P. Quinn, and V. R. R. Uppuluri. Some results on the best of paired comparisons. Applied Mathematics and Computation, 28(3):247–268, December 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300388901403.

Hempel:1988:NVA

[HT88] R. Hempel and C. Thompson. A note on the vectorization of algebraic multigrid algorithms. Applied Mathematics and Computation, 26(3):245–256, May 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388901221.

Ioakimidis:1985:NSA

[IA85] N. I. Ioakimidis and E. G. Anastasselou. A new, simple approach to the derivation of exact analytical formulae for the zeros of analytic functions. Applied Mathematics and Computation, 17(2): 123–127, September 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300385900220.

Itoh:1989:DLE

[IA89] Toshiaki Itoh and Kanji Abe. Discrete Lagrange’s equations and canonical equations based on the principle of least action. Applied Mathematics and Computation, 29(2):161–183, January 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300389901008.

Ierley:1986:SMS

[Ier86] Glenn R. Ierley. Spectral methods for the solution of nonlinear boundary value problems, a case study. Applied Mathematics and Computation, 20(1– 2):21, September 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386901190.

Ioakimidis:1988:AQR

[Ioa88] N. I. Ioakimidis. Application of quadrature rules to the determination of plane equipotential lines and other curves defined by Harmonic functions. Applied Mathematics and Computation, 27(2):147–154, August 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900252.

Ioakimidis:1989:NCQ

[IP89] N. I. Ioakimidis and K. E. Papadakis. A new class of quite elementary closed-form integral formulae for roots of nonlinear equations. Applied Mathematics and Computation, 29(2): 185–196, January 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038990101X.

Jameson:1985:PAA

[Jam85] Donald A. Jameson. A priori analysis of allowable interval between measurements as a test of model validity. Applied Mathematics and Computation, 17(2): 93–105, September 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300385900207.

Jameson:1986:SIM

[Jam86] Donald A. Jameson. Sampling intensity for monitoring of environmental systems. Applied Mathematics and Computation, 18(1):71–76, January 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300386900287.

Jameson:1988:IWV

[Jam88]  Donald A. Jameson.  Intro-

duction: workshop on vector and parallel processing in landscape dynamics. Applied Mathematics and Computation, 27(1):1–2, July 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630038890094X.

Jolayemi:1989:MCC

[JB89a] Joel K. Jolayemi and Julio N. Berrettoni. Multivariate control charts: an optimization approach to the effective use and measurement of performance. Applied Mathematics and Computation, 32(1):1–15, July 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038990043X.

Jolayemi:1989:ODM

[JB89b] Joel K. Jolayemi and Julio N.

Berrettoni. An optimal design of multivariate control charts in the presence of multiple assignable causes. Applied Mathematics and Computation, 32(1):17–33, July 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300389900441.

Jerrell:1988:RSS

[Jer88] Max E. Jerrell. A random search strategy for function optimization. Applied Mathematics and Computation, 28(3): 223–229, December 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388901385.

Jordan:1986:NSN

[JPS86] Kirk E. Jordan, George C. Papanicolaou, and R. Spigler. On the numerical solution of a nonlinear stochastic Helmholtz equation with a multigrid preconditioner. Applied Mathematics and Computation, 19(1– 4):145–157, July 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300386901013.

Kadalbajoo:1986:NSS

[Kad86] Mohan K. Kadalbajoo. Numerical solution of singular integral equations via initial value methods. Applied Mathematics and Computation, 18(3): 211–222, April 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300386900159.

Kamowitz:1988:MAS

[Kam88] David Kamowitz. Multigrid applied to singular perturbation problems. Applied Mathematics and Computation, 25(2): 145–174, January 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388901129.

Kapitza:1988:TIF

[Kap88] H. Kapitza. Truncated incomplete factorizations for conjugategradient methods in two and three dimensions. Applied Mathematics and Computation, 28 (1):73–87, October 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0096300388901348.

Kadalbajoo:1985:FDM

[KB85] Mohan K. Kadalbajoo and K. K.

Bharadwaj. On fast direct methods for solving elliptic equations over nonrectangular regions. Applied Mathematics and Computation, 16(1):1–17, January 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300385900074.

Kulev:1989:GSS

[KB89] G. K. Kulev and D. D. Bainov.

Global stability of sets for systems with impulses. Applied Mathematics and Computation, 29(3):255–270, February 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900179.

Keesman:1989:DPS

[Kee89] Karel Keesman. On the dominance of parameters in structural models of ill-defined systems. Applied Mathematics and Computation, 30(2):133–147, March 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389901471.

Kuttler:1987:SPH

[KH87] K. L. Kuttler and D. L. Hicks.

Some progress on the hydrocode convergence problem. Applied Mathematics and Computation, 23(3):211–233, September 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900798.

Kagiwada:1987:FIE

[KKS87] Harriet H. Kagiwada, Robert E. Kalaba, and Melvin R. Scott. The b and h functions for integral equations with displacement kernels: a computational method and an application to radiative transfer. Applied Mathematics and Computation, 23 (2):93–101, August 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900324.

Kuo:1989:TCF

[KL89] C.-C. Jay Kuo and Bernard C. Levy. Two-color Fourier analysis of the multigrid method with redblack Gauss–Seidel smoothing. Applied Mathematics and Computation, 29(1):69–87, January 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300389900404.

Kolp:1986:EMA

[KM86] O. Kolp and H. Mierendorff. Efficient multigrid algorithms for locally constrained parallel systems. Applied Mathematics and Computation, 19(1– 4):169–200, July 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386901037.

Kocvara:1987:MMT

[KM87] Michal Kočvara and Jan Mandel. A multigrid method for three-dimensional elasticity and algebraic convergence estimates. Applied Mathematics and Computation, 23(2):121–135, August 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300387900348.

Karplus:1986:CDS

[KN86] Kevin Karplus and Alexandru Nicolau. A compiler-driven supercomputer. Applied Mathematics and Computation, 20(1–2): 95–110, September 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386901281.

Kamowitz:1985:SSM

[KP85] David Kamowitz and Seymour V.

Parter. A study of some multigrid ideas. Applied Mathematics and Computation, 17(2):153– 184, September 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300385900256.

Kiviet:1986:TSM

[KP86] Jan F. Kiviet and Garry D. A. Phillips. Testing strategies for model specification. Applied Mathematics and Computation, 20(3–4):237–269, November 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630038690007X.

Kalaba:1987:ANM

[KPT87] Robert Kalaba, Thomas Plum, and Leigh Tesfatsion. Automation of nested matrix and derivative operations. Applied Mathematics and Computation, 23(3): 243–268, September 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900816.

Kadalbajoo:1987:AMN

[KR87a] Mohan K. Kadalbajoo and Y. N. Reddy. Approximate method for the numerical solution of singular perturbation problems. Applied Mathematics and Computation, 21(3):185–199, February 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900014.

Kadalbajoo:1987:NSS

[KR87b] Mohan K. Kadalbajoo and Y. N.

Reddy. Numerical solution of singular perturbation problems via deviating arguments. Applied Mathematics and Computation, 21(3):221–232, February 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900038.

Kadalbajoo:1987:NTS

[KR87c] Mohan K. Kadalbajoo and Y. N.

Reddy. Numerical treatment of singularly perturbed two point boundary value problems. Applied Mathematics and Computation, 21(2):93–110, February 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300387900208.

Kadalbajoo:1989:ANA

[KR89] Mohan K. Kadalbajoo and Y. N.

Reddy. Asymptotic and numerical analysis of singular perturbation problems: a survey. Applied Mathematics and Computation, 30(3):223–259, April 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900544.

Kunz:1986:CCS

[Kun86] A. Barry Kunz. Computational considerations in solving the many electron problem for molecular systems. Applied Mathematics and Computation, 20(1– 2):23–33, September 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386901207.

Kettler:1986:AMM

[KW86] R. Kettler and P. Wesseling. Aspects of multigrid methods for problems in three dimensions. Applied Mathematics and Computation, 19(1–4):159–168, July 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300386901025.

Lamberti:1986:SHH

[Lam86] Lamberto Lamberti. Solutions to the Hodgkin–Huxley equations. Applied Mathematics and Computation, 18(1):43–70, January 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300386900275.

Langerholc:1989:VDH

[Lan89] John Langerholc. Volumes of diced hyperspheres: resumming the Tam–Zardecki formula. Applied Mathematics and Computation, 30(1):1–18, March 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0096300389900623.

Larsen:1986:CFP

[Lar86] Jon T. Larsen. Computations in fusion physics. Applied Mathematics and Computation, 20(1– 2):111–142, September 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386901293.

Lether:1985:EAC

[LCE85] Frank G. Lether, David M. Cline, and Otis Evans. An error analysis for the calculation of semiintegrals and semiderivatives by the RL algorithm. Applied Mathematics and Computation, 17(1):45–67, July 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300385900189.

Lether:1985:ACR

[LE85] Frank G. Lether and Otis M.

Evans. An algorithm for the computation of the reversible Randles–Sevcik function in electrochemistry. Applied Mathematics and Computation, 16 (3):253–264, April 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300385900323.

Lessard:1989:MRA

[Les89] Sabin Lessard. The mathematics of resource allocation in population Biology. I. Interspecific competition. Applied Mathematics and Computation, 32(2– 3):207–226, August 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300389900945.

Lockhart:1986:PLL

[LH86] Deborah F. Lockhart and Darrell L. Hicks. Preface and list of lecturers. Applied Mathematics and Computation, 20(1–2): v–vii, September 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300386901165.

Lakshmikantham:1989:SID

[LL89] V. Lakshmikantham and Xinzhi Liu. Stability for impulsive differential systems in terms of two measures. Applied Mathematics and Computation, 29 (1):89–98, January 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300389900416.

Lubeck:1986:PET

[LMM86] Olaf M. Lubeck, James W. Moore, and Raul Mendez. A performance evaluation of three supercomputers, Fujitsu VP-200, Hitachi S810/20 and the Cray X-MP/24. Applied Mathematics and Computation, 20(1–2): 143, September 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038690130X.

Longman:1987:CMS

[Lon87a] I. M. Longman. Complementary moments and the summation of series. Applied Mathematics and Computation, 24(2): 139–151, November 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900658.

Longman:1987:ICR

[Lon87b] I. M. Longman. Increasing the convergence rate of series. Applied Mathematics and Computation, 24(1):77–89, October 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900312.

Longman:1987:SFC

[Lon87c] I. M. Longman. The summation of Fourier, Chebyshev, and Legendre series. Applied Mathematics and Computation, 23(1):61–70, July 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300387900567.

Lowther:1986:AMP

[Low86] John L. Lowther. AI machines and parallelism. Applied Mathematics and Computation, 20(1– 2):35, September 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300386901219.

Liu:1989:ASI

[LP89] Xinzhi Liu and Ratnam Pirapakaran. Asymptotic stability for impulsive differential systems. Applied Mathematics and Computation, 30(2):149–163, March 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300389901483.

Ladde:1985:DSM

[LR85] G. S. Ladde and S. G. Rajalakshmi. Diagonalization and stability of multi-time-scale singularly perturbed linear systems. Applied Mathematics and Computation, 16(2):115–140, February 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300385900037.

Lakshmikantham:1987:MUL

[LS87] V. Lakshmikantham and S. Sivasundaram. The method of upper and lower solutions and interval method for first order differential equations. Applied Mathematics and Computation, 23(1):1–5, July 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038790052X.

Lakshmikantham:1987:MMM

[LV87] V. Lakshmikantham and A. S.

Vatsala. Method of mixed monotony for nonlinear equations with a singular linear part. Applied Mathematics and Computation, 23(3):235–241, September 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300387900804.

Mandel:1986:MCN

[Man86] Jan Mandel. Multigrid convergence for nonsymmetric, indefinite variational problems and one smoothing step. Applied Mathematics and Computation, 19(1–4):201–216, July 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386901049.

Mandel:1988:ASM

[Man88] Jan Mandel. Algebraic study of multigrid methods for symmetric, definite problems. Applied Mathematics and Computation, 25(1):39–56, January 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630038890063X.

Matarazzo:1986:MAP

[Mat86]  Benedetto Matarazzo.  Mul-

ticriterion analysis of preferences by means of pairwise actions and criterion comparisons (MAPPACC). Applied Mathematics and Computation, 18(2): 119–141, February 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300386900202.

Mishev:1988:OSP

[MB88] D. P. Mishev and D. D. Bainov. Oscillation of the solutions of parabolic differential equations of neutral type. Applied Mathematics and Computation, 28(2): 97–111, November 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388900896.

McCormick:1986:P

[McC86] Steve McCormick. Preface. Applied Mathematics and Computation, 19(1–4):1, July 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300386900937.

Mitwasi:1989:AMD

[MCD89] Mousa Mitwasi, Ann-Marie Connolly, and Lucien Duckstein. Applying multicriteria decision making techniques for planning machining operations. Applied Mathematics and Computation, 29(3):197–218, February 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900131.

Meinhardt:1989:TCR

[Mei89] Hans Meinhardt. Tailoring and coupling of reaction-diffusion systems to obtain reproducible complex pattern formation during development of the higher organisms. Applied Mathematics and Computation, 32(2–3): 103–135, August 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900908.

Melson:1986:VMA

[Mel86] N. Duane Melson. Vectorizable multigrid algorithms for transonic-flow calculations. Applied Mathematics and Computation, 19(1–4):217–238, July 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386901050.

Merriam:1986:ADF

[Mer86] Marshal L. Merriam. Application of data flow concepts to a multigrid solver for the Euler equations. Applied Mathematics and Computation, 19 (1–4):239–263, July 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300386901062.

McGrath:1986:PAC

[MHL86] Joseph F. McGrath, Darrell L. Hicks, and Lorie M. Liebrock. Parallel algorithms for computational continuum dynamics. Applied Mathematics and Computation, 20(1–2):145–173, September 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300386901311.

Mikhail:1987:VSM

[Mik87] Mikhail Naguib Mikhail. On the validity and stability of the method of lines for the solution of partial differential equations. Applied Mathematics and Computation, 22(2–3):89–98, May 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900385.

Miller:1987:FRE

[Mil87] John D. Miller. Finding roots of equations in given neighborhoods. Applied Mathematics and Computation, 23(3):185– 192, September 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900774.

Miller:1988:FZA

[Mil88a] John Miller. Finding the zeros of an analytic function. Applied Mathematics and Computation, 27(3):243–249, September 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900057.

Milne:1988:MFG

[Mil88b] Bruce T. Milne. Measuring the fractal geometry of landscapes. Applied Mathematics and Computation, 27(1):67–79, July 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900999.

Mittelmann:1986:MCT

[Mit86] H. D. Mittelmann. Multilevel continuation techniques for nonlinear boundary value problems with parameter dependence. Applied Mathematics and Computation, 19(1–4):265–282, July 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386901074.

Miketinac:1987:APV

[MKJ87] M. J. Miketinac, A. W. Kelm, and D. S. Jayas. Airflow patterns for various approximations to velocity — pressure-gradient data. Applied Mathematics and Computation, 22(1):45–63, April 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630038790049X.

Miel:1985:CNL

[MM85] George Miel and Rose Mooney. On the condition number of Lagrangian numerical differentiation. Applied Mathematics and Computation, 16(3): 241–252, April 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300385900311.

Maitre:1987:MMS

[MM87] Jean-François Maitre and François Musy. Multigrid methods for symmetric variational problems: a general theory and convergence estimates for usual smoothers. Applied Mathematics and Computation, 21(1):21–43, January 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300387900075.

Martin:1989:NSP

[MMP89a] C. Martin, J. Miller, and K. Pearce. Numerical solution of positive sum exponential equations. Applied Mathematics and Computation, 34(2 (part 2)): 89–93, November 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900088.

Martin:1989:PIC

[MMP89b] C. Martin, J. Miller, and K. Pearce. Parameter identification by continuation methods. Applied Mathematics and Computation, 34(1):17–27, November 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300389900039.

Morgan:1986:HSP

[Mor86a] Alexander P. Morgan. A homotopy for solving polynomial systems. Applied Mathematics and Computation, 18(1): 87–92, January 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386900305.

Morgan:1986:TAS

[Mor86b] Alexander P. Morgan. A transformation to avoid solutions at infinity for polynomial systems. Applied Mathematics and Computation, 18(1):77–86, January 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300386900299.

Mosetti:1988:OSM

[Mos88a] R. Mosetti. Optimal strategies for monitoring a variable subjected to random changes. Applied Mathematics and Computation, 25(2):137–143, January 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300388901117.

Mosetti:1988:OPB

[Mos88b] Renzo Mosetti. Optimal policies in a Bioeconomic model of eutrophication. Applied Mathematics and Computation, 26 (3):257–265, May 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388901233.

McAleer:1986:SIN

[MP86] Michael McAleer and M. Hashem Pesaran. Statistical inference in non-nested econometric models. Applied Mathematics and Computation, 20(3–4): 271–311, November 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300386900081.

Muratori:1989:DSH

[MR89] Simona Muratori and Sergio Rinaldi. A dynamical system with Hopf bifurcations and catastrophes. Applied Mathematics and Computation, 29 (1):1–15, January 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900362.

Morgan:1987:CAS

[MS87a] Alexander Morgan and Andrew Sommese. Computing all solutions to polynomial systems using homotopy continuation. Applied Mathematics and Computation, 24(2):115–138, November 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900646.

Morgan:1987:HSG

[MS87b] Alexander Morgan and Andrew Sommese. A homotopy for solving general polynomial systems that respects m-homogeneous structures. Applied Mathematics and Computation, 24(2):101– 113, November 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900634.

Medhin:1988:MMM

[MS88] Negash G. Medhin and M. Sambandham. Mixed min-max optimization problem with restrictions. Applied Mathematics and Computation, 26(2):

151–167, May 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388900483.

Morgan:1989:CPP

[MS89] Alexander P. Morgan and Andrew J. Sommese. Coefficientparameter polynomial continuation. Applied Mathematics and Computation, 29(2): 123–160, January 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300389900994.

Muraskin:1989:SDL

[Mur89a] M. Muraskin. Study of different lattice solutions in aesthetic field theory. Applied Mathematics and Computation, 30 (1):73–98, March 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300389900659.

Muraskin:1989:UCR

[Mur89b] M. Muraskin. Use of commutation relations in no integrability aesthetic field theory. Applied Mathematics and Computation, 29(3):271–292, February 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300389900180.

Monsi:1988:IVS

[MW88] M. Monsi and M. A. Wolfe. Interval versions of some procedures for the simultaneous estimation of complex polynomial zeros. Applied Mathematics and Computation, 28(3):191–209, December 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300388901361.

Newbold:1986:TCU

[NH86] Paul Newbold and Steven M. Hotopp. Testing causality using efficiently parametrized vector ARMA models. Applied Mathematics and Computation, 20(3– 4):329–348, November 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038690010X.

Nicolau:1986:DES

[Nic86] Alexandru Nicolau. A development environment for scientific parallel programs. Applied Mathematics and Computation, 20(1– 2):175–183, September 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386901323.

Nieto:1985:SSS

[Nie85] Juan J. Nieto. On the structure of the solution set for first order differential equations. Applied Mathematics and Computation, 16(3):177–187, April 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300385900268.

Neta:1987:AMN

[NN87] Beny Neta and Paul Nelson. An adaptive method for the numerical solution of Fredholm integral equations of the second kind. I. regular kernels. Applied Mathematics and Computation, 21(2): 171–184, February 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900257.

Nowak:1989:GDA

[NS89] Martin Nowak and Karl Sigmund. Game-dynamical aspects of the prisoner’s dilemma. Applied Mathematics and Computation, 30(3):191–213, April 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900520.

Olson:1986:PST

[OM86] Robert Olson and Steve McGrogan. Parallelizing SPICE in a timesharing environment. Applied Mathematics and Computation, 20(1–2):185–195, September 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300386901335.

Omari:1986:MMC

[Oma86] Pierpaolo Omari. A monotone method for constructing extremal solutions of second order scalar boundary value problems. Applied Mathematics and Computation, 18(3):257–275, April 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386900184.

Ottenstein:1986:NRB

[Ott86] Karl J. Ottenstein. A note on the relationship between the expression and the automatic detection of parallelism in programs. Applied Mathematics and Computation, 20(1–2):37–40, September 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300386901220.

Palm:1986:SEM

[Pal86] F. C. Palm. Structural econometric modeling and time series analysis. Applied Mathematics and Computation, 20(3–4): 349–364, November 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386900111.

Pandian:1988:WDS

[Pan88] M. C. Pandian. A weighted difference scheme and monotone iterative methods for quasilinear boundary value problems. Applied Mathematics and Computation, 25(3):187–205, February 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300388900719.

Parter:1985:NCM

[Par85] Seymour V. Parter. A note on convergence of the multigrid V -cycle. Applied Mathematics and Computation, 17(2): 137–151, September 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300385900244.

Parter:1987:RMC

[Par87] Seymour V. Parter. Remarks on multigrid convergence theorems. Applied Mathematics and Computation, 23(2):103–120, August 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300387900336.

Pardalos:1989:PSA

[Par89] P. M. Pardalos. Parallel search algorithms in global optimization. Applied Mathematics and Computation, 29(3):219– 229, February 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900143.

Patterson:1987:REI

[Pat87] Richard L. Patterson. Regression estimates of inputs to an M(t)/G/∞ service system. Applied Mathematics and Computation, 24(1):47–63, October 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900294.

Patterson:1989:RES

[Pat89]  Richard L. Patterson.  Re-

gression estimates of stochastic compartmental parameters using net balance and cumulative flow data. Applied Mathematics and Computation, 29(3): 245–254, February 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900167.

Peters:1988:PES

[Pet88] Randall D. Peters. Period estimate for the simple pendulum using optimization theory. Applied Mathematics and Computation, 28(3):179–190, December 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038890135X.

Pignatiello:1987:EMO

[Pig87] Joseph J. Pignatiello, Jr. Enumerating maximal-order abelian groups within a set of eligible elements. Applied Mathematics and Computation, 22 (2–3):99–114, May 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900397.

Luo:1987:EVAa

[pLR87a] Zi ping Luo and Herschel Rabitz. Expected value analysis: Experimental considerations and convergence properties. Applied Mathematics and Computation, 23(1):25–47, July 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900543.

Luo:1987:EVAb

[pLR87b] Zi ping Luo and Herschel Rabitz. Expected value analysis of multiparameter systems. Applied Mathematics and Computation, 24(2):153–181, November 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630038790066X.

Patterson:1988:SAP

[PM88] Richard L. Patterson and Zhen Qui Ma. Sensitivity analyses of parameters of a M(t)/G/∞ stochastic service system. Applied Mathematics and Computation, 26(2):169–177, May 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900495.

Patterson:1989:SDC

[PM89] Richard L. Patterson and Zhenkui Ma. Statistical distributions of compartmentalized populations governed by continuous time, discrete state semi-Markov processes. Applied Mathematics and Computation, 30(1): 49–71, March 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900647.

Poplawski:1986:PCA

[Pop86] David A. Poplawski. Parallel computer architectures. Applied Mathematics and Computation, 20(1–2):41–51, September 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386901232.

Ravn:1987:FMP

[Rav87] Hans F. Ravn. A forward maximum principle algorithm with decision horizon results. Applied Mathematics and Computation, 24(1):65–75, October 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900300.

Reed:1987:OPM

[Ree87] William J. Reed. Optimal preventive maintenance, protection, and replacement of a revenueearning asset. Applied Mathematics and Computation, 24(3): 241–261, December 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900877.

Rand:1987:DDE

[RK87] R. H. Rand and W. L. Keith. Determinacy of degenerate equilibria with linear part x′ = y, y′ = 0 using MACSYMA. Applied Mathematics and Computation, 21(1):1–19, January 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900063.

Rogers:1986:SPI

[RMR86] C. Rogers, C. McInnis, and H. Rasmussen. On a Stefan problem involving a spherical region: application of a Bergmantype expansion. Applied Mathematics and Computation, 18(2): 143–154, February 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300386900214.

Roache:1986:ASE

[RS86] Patrick J. Roache and Stanly Steinberg. Application of a single-equation MG-FAS solver to elliptic grid generation equations (subgrid and supergrid coefficient generation). Applied Mathematics and Computation, 19(1–4):283–292, July 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386901086.

Rao:1989:ABS

[RSS89] M. Rama Mohana Rao, S. Sathananthan, and S. Sivasundaram. Asymptotic behavior of solutions of impulsive integrodifferential systems. Applied Mathematics and Computation, 34(3): 195–211, December 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389901045.

Ruge:1986:APE

[Rug86] J. Ruge. AMG for problems of elasticity. Applied Mathematics and Computation, 19(1– 4):293–309, July 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386901098.

Said:1988:IP

[Sai88] H. B. Said. C^{1} interpolation on a polygon. Applied Mathematics and Computation, 27(3): 217–229, September 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900033.

Sasagawa:1985:ASO

[Sas85]  T. Sasagawa.  Analysis of

the second order matrix Riccati equations. Applied Mathematics and Computation, 16(2): 141–152, February 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300385900049.

Sasagawa:1988:ETD

[Sas88] T. Sasagawa. Existence theorem for the difference Riccati equations. Applied Mathematics and Computation, 26(2):89–103, May 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300388900446.

Scandrett:1989:CSI

[Sca89] C. Scandrett. Comparison of several iterative techniques in the solution of symmetric banded equations on a two-pipe cyber 205. Applied Mathematics and Computation, 34(2 (part 2)): 95–112, November 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/009630038990009X.

Shabahang:1987:MWS

[Sha87] R. Shabahang. Mean wave solution: bounded medium — discrete scatterers. Applied Mathematics and Computation, 21(2): 111–123, February 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038790021X.

Shalaby:1989:WUL

[Sha89] Mohamed Ahmed Shalaby. Weighted and unweighted least-squares solutions of linear matrix equations. Applied Mathematics and Computation, 34(1): 61–71, November 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900064.

Shivaji:1989:RSC

[Shi89] Ratnasingham Shivaji. Remark on a sufficient condition for multiplicity in a model in chemical reactor theory. Applied Mathematics and Computation, 29 (2):99–105, January 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900970.

Shyy:1987:AGM

[Shy87] Wei Shyy. An adaptive grid method for Navier–Stokes flow computation. Applied Mathematics and Computation, 21(3): 201–219, February 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900026.

Swisshelm:1986:PCE

[SJK86] J. M. Swisshelm, G. M. Johnson, and S. P. Kumar. Parallel computation of Euler and Navier–Stokes flows. Applied Mathematics and Computation, 19(1–4):321–331, July 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300386901116.

Scott:1988:SEM

[SM88] Melvin R. Scott and Gary R. Montry. Some experiments in multitasking on an ELXSI system 6400. Applied Mathematics and Computation, 26 (2):135–150, May 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388900471.

Smith:1986:TET

[Smi86] Tony E. Smith. Two extension theorems for log-linear representations of positive probabilities.

Applied Mathematics and Computation, 18(3):239–255, April 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300386900172.

Sochacki:1988:ABC

[Soc88]  James Sochacki.  Absorbing

boundary conditions for the elastic wave equations. Applied Mathematics and Computation, 28(1):1–14, October 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388901300.

Sosnowska:1988:GSB

[Sos88] Honorata Sosnowska. Global stability in bounded sets: Application to a queueing model of a market in a shortage economy. Applied Mathematics and Computation, 27(2):155–165, August 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300388900264.

Shi:1988:OMS

[SR88] Shenghua Shi and Herschel Rabitz. An operator method for solving the determining equations for the group analysis of differential equations. Applied Mathematics and Computation, 25(2): 101–113, January 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388901087.

Schaffer:1986:MSM

[SS86] Steve Schaffer and Frank Stenger. Multigrid-sinc methods. Applied Mathematics and Computation, 19(1–4):311–319, July 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300386901104.

Szidarovszky:1987:MOM

[SS87] Ferenc Szidarovszky and Károly Szenteleki. A multiobjective optimization model for wine production. Applied Mathematics and Computation, 22(2– 3):255–275, May 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300387900452.

Sugisaka:1987:DLR

[Sug87] Masanori Sugisaka. The design of a linear regulator based on an integral-equation representation. Applied Mathematics and Computation, 23(2):163–169, August 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300387900361.

Saaty:1985:MBC

[SV85] Thomas L. Saaty and Luis G. Vargas. Modeling behavior in competition: the analytic hierarchy process. Applied Mathematics and Computation, 16 (1):49–92, January 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300385900098.

So:1989:NBV

[SW89] Joseph W.-H. So and Paul Waltman. A nonlinear boundary value problem arising from competition in the chemostat. Applied Mathematics and Computation, 32(2– 3):169–183, August 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900921.

Szidarovszky:1989:GST

[Szi89]  Ferenc Szidarovszky.  On a

general scheme in the theory of conflicts. Applied Mathematics and Computation, 34(1): 29–37, November 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900040.

Thole:1986:BSP

[TT86] Clemens-August Thole and Ulrich Trottenberg. Basic smoothing procedures for the multigrid treatment of elliptic 3D operators. Applied Mathematics and Computation, 19(1– 4):333–345, July 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300386901128.

Thole:1988:SNS

[TT88] Clemens-August Thole and Ulrich Trottenberg. A short note on standard parallel multigrid algorithms for 3D problems. Applied Mathematics and Computation, 27(2):101–115, August 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388900239.

Tennakone:1989:PIF

[TT89] K. Tennakone and D. Tantirigoda. Periodicity in the incidence of filariasis and the solar cycle. Applied Mathematics and Computation, 32(1):35–43, July 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300389900453.

Tucker:1988:BIC

[TTB88] Jerry H. Tucker, Moiez A. Tapia, and A. Wayne Bennett. Boolean integral calculus. Applied Mathematics and Computation, 26 (3):201–236, May 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300388901208.

Turner:1988:SSM

[Tur88] Monica Goigel Turner. A spatial simulation model of land use changes in a Piedmont county in Georgia. Applied Mathematics and Computation, 27 (1):39–51, July 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900975.

Tennakone:1989:CSE

[TW89] K. Tennakone and S. Wickremanayake. Computer study of the effect of noise on an epidemiological model. Applied Mathematics and Computation, 34(2 (part 2)):73–87, November 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900076.

Usmani:1988:NST

[UA88] Riaz A. Usmani and Ravi P.

Agarwal. On the numerical solution of two point discrete boundary value problems. Applied Mathematics and Computation, 25(3):247–264, February 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900756.

Udwadia:1987:RUDa

[Udw87a] F. E. Udwadia. Response of uncertain dynamic systems. I. Applied Mathematics and Computation, 22(2–3):115–150, May 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900403.

Udwadia:1987:RUDb

[Udw87b] F. E. Udwadia. Response of uncertain dynamic systems. II. Applied Mathematics and Computation, 22(2–3):151–187, May 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900415.

Ueno:1989:SAC

[Uen89] S. Ueno. Stochastic approach to the Cauchy system for the scattering function with hybrid reflector. Applied Mathematics and Computation, 29 (1):39–47, January 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900386.

Venkatachalam:1988:PTU

[VA88] P. A. Venkatachalam and S. Arumugam. Program termination using Z-transform theory. Applied Mathematics and Computation, 27(3):231–241, September 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900045.

Valyi:1988:SAD

[Vál88] István Vályi. Strict approximate duality in vector spaces. Applied Mathematics and Computation, 25(3):227–246, February 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900744.

Vincent:1989:ERS

[VB89] Thomas L. Vincent and Joel S.

Brown. The evolutionary response of systems to a changing environment. Applied Mathematics and Computation, 32(2– 3):185–206, August 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900933.

Vargas-Jarillo:1986:DMR

[VJ86] Cristobal Vargas-Jarillo. A discrete model for the recovery of oil from a reservoir. Applied Mathematics and Computation, 18(2): 93–117, February 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386900196.

Vazquez:1988:MCM

[VJ88] Luis Vázquez and Salvador Jiménez. Motion of a charge in a magnetic dipole field. I. Painlevé analysis and a conservative numerical scheme. Applied Mathematics and Computation, 25(3):

207–217, February 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388900720.

Watson:1989:SFD

[Wat89] Ray Watson. On the spread of a fatal disease. Applied Mathematics and Computation, 33(2): 123–129, September 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900866.

Wang:1988:AMQ

[WC88] H. Wang and A. J. Chapman. Application of the modified quasilinearization algorithm to the construction of the contour of turbine blades. Applied Mathematics and Computation, 25(2): 115–122, January 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300388901099.

Wexler:1987:AED

[Wex87] Anthony S. Wexler. Automatic evaluation of derivatives. Applied Mathematics and Computation, 24(1):19–46, October 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900282.

Wang:1985:DT

[WH85] Zhongde Wang and B. R. Hunt.

The discrete W transform. Applied Mathematics and Computation, 16(1):19–48, January 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300385900086.

Wang:1989:SCH

[WHF89] Muhong Wang, Keith W. Hipel, and Niall M. Fraser. Solution concepts in hypergames. Applied Mathematics and Computation, 34(3):147–171, December 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389901021.

White:1989:EHO

[Whi89] L. W. White. Estimation of higher-order damping terms in linear plate models. Applied Mathematics and Computation, 33(2):89–122, September 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900854.

Wexler:1988:IIS

[WKTM88] A. S. Wexler, R. E. Kalaba, L. S. Tesfatsion, and D. J. Marsh. An invariant-imbedding solution of general linear two-point boundary-value problems. Applied Mathematics and Computation, 26(3):237–244, May 1988.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038890121X.

Wolfe:1989:OIM

[Wol89] M. A. Wolfe. On order-interval methods for bounding zeros of order convex operators. Applied Mathematics and Computation, 32(1):45–59, July 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900465.

Yserentant:1986:HBG

[Yse86] Harry Yserentant. Hierarchical bases give conjugate gradient type methods a multigrid speed of convergence. Applied Mathematics and Computation, 19(1–4):347–358, July 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038690113X.

Zemanian:1988:FDA

[ZA88] A. H. Zemanian and Hyeong Keon An. Finite-difference analysis of borehole flows involving domain contractions around threedimensional anomalies. Applied Mathematics and Computation, 26(1):45–75, April 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0096300388900860.

Zakeri:1989:GSW

[Zak89] Gholam-Ali Zakeri. Geometric structure of 2D weak shock waves. Applied Mathematics and Computation, 33(3): 161–183, October 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900490.

Zamani:1988:BES

[Zam88] N. G. Zamani. Boundary element simulation of the cathodic protection system in a prototype ship. Applied Mathematics and Computation, 26(2):119–134, May 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630038890046X.

Zeldin:1987:TAL

[Zel87] Yevgeny Zeldin. Topological analysis of large networks. Applied Mathematics and Computation, 23(3):193–210, September 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300387900786.

Zemanian:1986:MMN

[Zem86] A. H. Zemanian. Multilevel marketing networks in less-developed countries. Applied Mathematics and Computation, 18(3):

185–210, April 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386900147.

Zang:1986:SMM

[ZH86] Thomas A. Zang and M. Yousuff Hussaini. On spectral multigrid methods for the time-dependent Navier–Stokes equations. Applied Mathematics and Computation, 19(1–4):359–372, July 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300386901141.

Zhang:1989:SRI

[Zha89] Wei-Bin Zhang. Short-run inventory oscillations in the Eckalbar disequilibrium macro model. Applied Mathematics and Computation, 33(1):53–67, September 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300389900593.

Zia:1989:EPT

[Zia89] Lee L. Zia. Estimation of parameters in two-dimensional transport equations. Applied Mathematics and Computation, 30 (1):19–47, March 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300389900635.

Zhu-quan:1987:NTL

[Zq87] Zang Zhu-quan. A note on the theorem of Lakshmikantham and Leela. Applied Mathematics and Computation, 23(3):269– 275, September 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300387900828.